OpenStudy (anonymous):

log base x - log base 6 =log base 15

OpenStudy (anonymous):

is the problem written correctly? base x or base ???

OpenStudy (anonymous):

Assuming that is log in base 10, raise 10 to both sides: 10^log(x) = 10^6 x = 1,000,000

OpenStudy (anonymous):

dpalnc, yes the question is written correctly.

OpenStudy (anonymous):

oh yeah....

OpenStudy (anonymous):

Rohangrr...the answer is 90...I just need to know how to get there.

OpenStudy (anonymous):

@ roh... base 10 yep.

OpenStudy (anonymous):

you can write that left side as logbase (x/6) then .... and do the same as the previous problem...

OpenStudy (anonymous):

so x/6 = 15

OpenStudy (anonymous):

oh ok, I see now...then to undo x/6..you multiply 6 by both sides and 15 * 6 =90!!!

OpenStudy (anonymous):

ye